# When Attention Sink Emerges in Language Models: An Empirical View

### v1

@@ 8 Future work
This work focuses on the sink token in the first position. Sun et al. (2024); Yu et al. (2024) showed that attention sink can also appear on certain word tokens, e.g., period and newline tokens. However, these sink words may vary in different LMs and typically have no fixed positions. In the future, we will extend the research scope to explore how these sink tokens are related to the pre-training. Additionally, we will also consider learnable KV biases in continual pre-training or instruction tuning.
Appendix A Detailed formulations of positional embedding
@@
In this section, we provide detailed formulations of positional embedding (PE) in LMs. PEs could be classified into two categories. NoPE, absolute positional embedding, and learnable positional embedding belong to the same category since they are added to the initial hidden states: $\bm{H}^{0}=\bm{X}\bm{W}_{E}+\bm{P}$. Here $\bm{P}=\left\{\bm{p}_{1}\textrm{,}\,\bm{p}_{2}\textrm{,}\,\cdots\textrm{,}\,\bm{p}_{T}\right\}\in\mathbb{R}^{T\times d}$. Meanwhile, the dot product between each query and key is computed as $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}^{T}\bm{k}_{j}$.
@@
Relative PE. The relative positional embeddings are adopted in T5 models (Raffel et al., 2020). A bias term is adopted for the dot product: $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{k}_{j}^{T}+g(i-j)$, where the definition for distance function $g(\cdot)$ is:
@@
ALiBi. Similarly, ALiBi (Press et al., 2021) also adds a bias term to the dot product: $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}^{T}\bm{k}_{j}+g(i-j)$, where $g(i-j)=-(i-j)\cdot m$. $m$ is a head-specific slope fixed:
@@
$\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\left(\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,-i}\right)\left(\bm{k}_{j}\bm{R}_{\Theta\textrm{,}\,-j}\right)^{T}=\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,-i}\bm{R}_{\Theta\textrm{,}\,j}\bm{k}_{j}^{T}=\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,j-i}\bm{k}_{j}^{T}\textrm{,}$ (8)
@@
Here $\omega_{i}=1/\textrm{10000}^{2(i-1)/d_{h}}$ and $d_{h}=d/H$ is the hidden dimension in each head. From the above definition, it is noted that the rotation matrix $\bm{R}_{\Theta,m}^{d_{h}}$ satisfies $\bm{R}_{\Theta,m}^{T}=\bm{R}_{\Theta,-m}$ and $\bm{R}_{\Theta,i}\bm{R}_{\Theta,j}=\bm{R}_{\Theta,i+j}$.
Appendix B Attention sink in open-sourced LMs
B.1 How positional embedding relates with attention sink
@@
Typically, the hidden states of the first token $\bm{h}_{1}^{l}$ in specific blocks have massive activations. Due to the above equality, all the repeated tokens have massive activations.
For LMs with absolute PE/learnable PE, the initial hidden states $\bm{h}_{t}^{0}=\bm{x}\bm{W}_{E}+\bm{p}_{t}$. Although the word embeddings are the same for repeated tokens, $\bm{p}_{t}$ for different token positions $t$ is different. Therefore, GPT2 models have no the above equality.
@@
$\bm{A}_{ti}^{l\textrm{,}\,h}=\frac{e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle}}{\sum_{j=1}^{t}e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h}\rangle}}=\frac{e^{\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{i}^{l\textrm{,}h}}^{T}}}{\sum_{j=1}^{t}e^{\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{j}^{l\textrm{,}h}}^{T}}}=\frac{e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{T}}}{te^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{T}}}=\frac{1}{t}\textrm{.}$ (18)
@@
$\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle=\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{i}^{l\textrm{,}h}}^{T}+g_{\textrm{rel}}(t-i)=\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{T}+g_{\textrm{rel}}(t-i)\textrm{,}$ (19)
@@
$\bm{A}_{t\textrm{,}\,i}^{l\textrm{,}\,h}=\frac{e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle}}{\sum_{j=1}^{t}e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h}\rangle}}=\frac{e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{T}+g_{\textrm{rel}}(t-i)}}{\sum_{j=1}^{t}e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{T}+g_{\textrm{rel}}(t-j)}}=\frac{e^{g_{\textrm{rel}}(t-i)}}{\sum_{j=1}^{t}e^{g_{\textrm{rel}}(t-j)}}\textrm{.}$ (20)
@@
$\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle=\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{i}^{l\textrm{,}h}}^{T}+g_{\textrm{alibi}}^{h}(t-i)=\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{T}+g_{\textrm{alibi}}^{h}(t-i)\textrm{,}$ (21)
@@
$\bm{A}_{t\textrm{,}\,i}^{l\textrm{,}\,h}=\frac{e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle}}{\sum_{j=1}^{t}e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h}\rangle}}=\frac{e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{T}+g_{\textrm{alibi}}^{h}(t-i)}}{\sum_{j=1}^{t}e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{T}+g_{\textrm{alibi}}^{h}(t-j)}}=\frac{e^{g_{\textrm{alibi}}^{h}(t-i)}}{\sum_{j=1}^{t}e^{g_{\textrm{alibi}}^{h}(t-j)}}\textrm{.}$ (22)
@@
$\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle$ $=\bm{q}_{t}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,i-t}{\bm{k}_{i}^{l\textrm{,}h}}^{T}$ (23)
$=\bm{q}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,i-t}{\bm{k}^{l\textrm{,}h}}^{T}$ (24)
$=\left\|\bm{q}^{l\textrm{,}h}\right\|\left\|\bm{k}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,t-i}\right\|\textrm{cos}\left(\frac{\bm{q}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,i-t}{\bm{k}^{l\textrm{,}h}}^{T}}{\left\|\bm{q}^{l\textrm{,}h}\right\|\left\|\bm{k}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,t-i}\right\|}\right)$ (25)
@@
$\bm{A}_{t\textrm{,}\,i}^{l\textrm{,}\,h}=\frac{e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle}}{\sum_{j=1}^{t}e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h}\rangle}}=\frac{e^{\bm{q}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,j-i}{\bm{k}^{l\textrm{,}h}}^{T}}}{\sum_{j=1}^{t}e^{\bm{q}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,j-i}{\bm{k}^{l\textrm{,}h}}^{T}}}=\frac{e^{\left\|\bm{q}^{l\textrm{,}h}\right\|\left\|\bm{k}^{l\textrm{,}h}\right\|\textrm{cos}(\beta_{t-i})}}{\sum_{j=1}^{t}e^{\left\|\bm{q}^{l\textrm{,}h}\right\|\left\|\bm{k}^{l\textrm{,}h}\right\|\textrm{cos}(\beta_{t-j})}}\textrm{.}$ (27)
@@
B.2 Attention sink under different data domains

### v2

@@ 8 Future work
This work focuses on the sink token in the first position. Sun et al. (2024); Yu et al. (2024) showed that attention sink can also appear on certain word tokens, e.g., period and newline tokens. However, these sink words may vary in different LMs and typically have no fixed positions. In the future, we will extend the research scope to explore how these sink tokens are related to the pre-training. Additionally, it remains unclear whether attention sink benefits LM downstream performance.
Appendix A Related Work
A.1 Attention sink phenomenon
Although the term attention sink was introduced by Xiao et al. (2023b), similar observations have been reported in prior studies. Among them, Zhai et al. (2023) demonstrated that attention weights in Transformer models (Vaswani et al., 2017) collapse into one-hot vectors, a phenomenon termed attention entropy collapse. This issue has been identified as attention logit growth by Wortsman et al. (2023); Dehghani et al. (2023). Furthermore, Bondarenko et al. (2023) observed the emergence of strong outliers in Transformer models, linked to attention heads that learn not to update residuals. Although not explicitly using the term attention sink, Han et al. (2024) highlighted the significance of the first few tokens in language models (LMs), noting their distinct representational space. In vision transformers (ViTs), Darcet et al. (2023) identified high-norm outlier tokens as artifacts in attention maps. Overall, the attention sink phenomenon is prevalent across Transformer models, including ViTs, encoder-only LMs, such as BERT (Devlin et al., 2019), and auto-regressive LMs. Notably, in auto-regressive LMs, attention sink consistently occurs on the first token in addition to tokens with limited semantic information in the middle context (Sun et al., 2024; Yu et al., 2024).
A.2 Attention sink and activation outliers
Cancedda (2024) found that early FFNs in LLaMA2 blast off the large norm of hidden states for the first token, leading to attention sink in later layers. This is referred to as massive activations (very few activations exhibit disproportionately large values) in Sun et al. (2024) and token-wise activation outlier in outlier literature (Lin et al., 2024). Similar observations were made by Bondarenko et al. (2023). In contrast, channel-wise activation outlier (Dettmers et al., 2022; Xiao et al., 2023a) refer to that outliers appear in specific channels across all token positions. A concurrent study (Kaul et al., 2024) found that these two types of outliers are distinct and require different mitigation strategies. Despite the high magnitude in hidden states, Bondarenko et al. (2023); Guo et al. (2024b) observed the values associated with sink tokens are typically smaller than those of other tokens.
A.3 Understanding and mitigate attention sink
To understand the attention sink phenomenon in general Transformer models, Bondarenko et al. (2023) hypothesized that the interplay among softmax, residual connections, and LayerNorm encourages models to learn not to update residuals. In auto-regressive LMs, Sun et al. (2024) attributed attention sink at the first position to implicit biases in keys and values. A concurrent work by Guo et al. (2024a) provided empirical and theoretical analyses of attention sink in very simple Transformer LMs on the Bigram-Backcopy (BB) task, identifying a mutual reinforcement mechanism as the underlying cause. Their findings on the BB task indicated that (i) replacing softmax with ReLU removes attention sink; (ii) switching the optimizer Adam to SGD eliminates massive activations but attention sink remains.
Besides the above literature, Xiao et al. (2023b); Kaul et al. (2024) found that replacing softmax with softmax-off-by-one (Miller, 2023) alleviates attention sink at the first position, though Xiao et al. (2023b) noted its potential emergence at other initial token positions. Yin et al. (2024) proposed refining the casual mask with pseudo-attention values on the “future tokens”, which could reduce attention sink. Though not explicitly targeting attention sink, methods such as $\sigma$Reparam (Zhai et al., 2023) and qk-norm (Dehghani et al., 2023) were introduced to address attention entropy collapse. To mitigate activation outliers, He et al. (2024) developed an outlier-protected block and also highlighted the role of optimization in mitigating the outliers. Besides, Bondarenko et al. (2023) proposed clipped softmax and gated attention, while Kaul et al. (2024) introduced OrthoAdam to reduce activation outliers.
A.4 Applications of attention sink
Sink tokens in LMs are functionally important and absorb significant attention. This enables computational savings by prioritizing initial and recent tokens over middle-context tokens in attention calculations. Token-wise activation outliers pose challenges for quantization, necessitating specialized handling to facilitate quantization processes. These insights have motivated various applications, including streaming/long context generation (Xiao et al., 2023b; Han et al., 2024; Yang et al., 2024), KV cache optimization (Ge et al., 2023; Wan et al., 2024; Wu & Tu, 2024), efficient inference (Zhang et al., 2024b; Chen et al., 2024), model quantization (Liu et al., 2024b; Huang et al., 2024), and others.
Appendix B Detailed formulations of positional embedding
@@
In this section, we provide detailed formulations of positional embedding (PE) in LMs. PEs could be classified into two categories. NoPE, absolute positional embedding, and learnable positional embedding belong to the same category since they are added to the initial hidden states: $\bm{H}^{0}=\bm{X}\bm{W}_{E}+\bm{P}$. Here $\bm{P}=\left\{\bm{p}_{1}\textrm{,}\,\bm{p}_{2}\textrm{,}\,\cdots\textrm{,}\,\bm{p}_{T}\right\}\in\mathbb{R}^{T\times d}$. Meanwhile, the dot product between each query and key is computed as $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{k}_{j}^{\top}$.
@@
Relative PE. The relative positional embeddings are adopted in T5 models (Raffel et al., 2020). A bias term is adopted for the dot product: $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{k}_{j}^{\top}+g(i-j)$, where the definition for distance function $g(\cdot)$ is:
@@
ALiBi. Similarly, ALiBi (Press et al., 2021) also adds a bias term to the dot product: $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\bm{q}_{i}\bm{k}_{j}^{\top}+g(i-j)$, where $g(i-j)=-(i-j)\cdot m$. $m$ is a head-specific slope fixed:
@@
$\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle=\left(\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,-i}\right)\left(\bm{k}_{j}\bm{R}_{\Theta\textrm{,}\,-j}\right)^{\top}=\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,-i}\bm{R}_{\Theta\textrm{,}\,j}\bm{k}_{j}^{\top}=\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,j-i}\bm{k}_{j}^{\top}\textrm{,}$ (8)
@@
Here $\omega_{i}=1/\textrm{10000}^{2(i-1)/d_{h}}$ and $d_{h}=d/H$ is the hidden dimension in each head. From the above definition, it is noted that the rotation matrix $\bm{R}_{\Theta,m}^{d_{h}}$ satisfies $\bm{R}_{\Theta,m}^{\top}=\bm{R}_{\Theta,-m}$ and $\bm{R}_{\Theta,i}\bm{R}_{\Theta,j}=\bm{R}_{\Theta,i+j}$.
Appendix C Attention sink in open-sourced LMs
C.1 How positional embedding relates to attention sink
@@
Typically, the hidden states of the first token $\bm{h}_{1}^{l}$ in specific blocks have massive activations. Due to the above equality, all the repeated tokens have massive activations. Furthermore, we could derive the closed form or upper bounds for attention scores under the repeated token sequence.
@@
$\bm{A}_{ti}^{l\textrm{,}\,h}=\frac{e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle}}{\sum_{j=1}^{t}e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h}\rangle}}=\frac{e^{\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{i}^{l\textrm{,}h}}^{\top}}}{\sum_{j=1}^{t}e^{\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{j}^{l\textrm{,}h}}^{\top}}}=\frac{e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{\top}}}{te^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{\top}}}=\frac{1}{t}\textrm{.}$ (18)
@@
$\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle=\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{i}^{l\textrm{,}h}}^{\top}+g_{\textrm{rel}}(t-i)=\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{\top}+g_{\textrm{rel}}(t-i)\textrm{,}$ (19)
@@
$\bm{A}_{t\textrm{,}\,i}^{l\textrm{,}\,h}=\frac{e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle}}{\sum_{j=1}^{t}e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h}\rangle}}=\frac{e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{\top}+g_{\textrm{rel}}(t-i)}}{\sum_{j=1}^{t}e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{\top}+g_{\textrm{rel}}(t-j)}}=\frac{e^{g_{\textrm{rel}}(t-i)}}{\sum_{j=1}^{t}e^{g_{\textrm{rel}}(t-j)}}\textrm{.}$ (20)
@@
$\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle=\bm{q}_{t}^{l\textrm{,}h}{\bm{k}_{i}^{l\textrm{,}h}}^{\top}+g_{\textrm{alibi}}^{h}(t-i)=\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{\top}+g_{\textrm{alibi}}^{h}(t-i)\textrm{,}$ (21)
@@
$\bm{A}_{t\textrm{,}\,i}^{l\textrm{,}\,h}=\frac{e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle}}{\sum_{j=1}^{t}e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h}\rangle}}=\frac{e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{\top}+g_{\textrm{alibi}}^{h}(t-i)}}{\sum_{j=1}^{t}e^{\bm{q}^{l\textrm{,}h}{\bm{k}^{l\textrm{,}h}}^{\top}+g_{\textrm{alibi}}^{h}(t-j)}}=\frac{e^{g_{\textrm{alibi}}^{h}(t-i)}}{\sum_{j=1}^{t}e^{g_{\textrm{alibi}}^{h}(t-j)}}\textrm{.}$ (22)
@@
$\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle$ $=\bm{q}_{t}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,i-t}{\bm{k}_{i}^{l\textrm{,}h}}^{\top}$ (23)
$=\bm{q}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,i-t}{\bm{k}^{l\textrm{,}h}}^{\top}$ (24)
$=\left\|\bm{q}^{l\textrm{,}h}\right\|\left\|\bm{k}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,t-i}\right\|\textrm{cos}\left(\frac{\bm{q}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,i-t}{\bm{k}^{l\textrm{,}h}}^{\top}}{\left\|\bm{q}^{l\textrm{,}h}\right\|\left\|\bm{k}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,t-i}\right\|}\right)$ (25)
@@
$\bm{A}_{t\textrm{,}\,i}^{l\textrm{,}\,h}=\frac{e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{i}^{l\textrm{,}h}\rangle}}{\sum_{j=1}^{t}e^{\langle\bm{q}_{t}^{l\textrm{,}h}\textrm{,}\,\bm{k}_{j}^{l\textrm{,}h}\rangle}}=\frac{e^{\bm{q}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,j-i}{\bm{k}^{l\textrm{,}h}}^{\top}}}{\sum_{j=1}^{t}e^{\bm{q}^{l\textrm{,}h}\bm{R}_{\Theta\textrm{,}\,j-i}{\bm{k}^{l\textrm{,}h}}^{\top}}}=\frac{e^{\left\|\bm{q}^{l\textrm{,}h}\right\|\left\|\bm{k}^{l\textrm{,}h}\right\|\textrm{cos}(\beta_{t-i})}}{\sum_{j=1}^{t}e^{\left\|\bm{q}^{l\textrm{,}h}\right\|\left\|\bm{k}^{l\textrm{,}h}\right\|\textrm{cos}(\beta_{t-j})}}\textrm{.}$ (27)
@@
For LMs with absolute PE/learnable PE, the initial hidden states $\bm{h}_{t}^{0}=\bm{x}\bm{W}_{E}+\bm{p}_{t}$. Although the word embeddings are the same for repeated tokens, $\bm{p}_{t}$ for different token positions $t$ is different. Therefore, GPT2 models have no the above equality. From Table 1(Left), GPT2-XL still allocates significant attention to the first token even with repeated tokens, which motivates us to explore whether attention sink is related to these learned positional embedding vectors $\bm{p}_{1:T}$ after LM pre-training.
Therefore, we conduct two experiments on GPT2-XL. Firstly, we replace the first positional embedding vector $\bm{p}_{1}$ with other vectors $\bm{p}_{t\neq 1}$. In Table 7, we find that the amplitude of attention sink on the first token is significantly reduced. Then we also consider swapping the first positional embedding vector $\bm{p}_{1}$ with another position $\bm{p}_{t\neq 1}$. Consequently, the $t$-th token becomes the new sink token. Therefore, attention sink in GPT2-XL is strongly attached to the first positional embedding vector $\bm{p}_{1}$.
Table 7: In GPT2-XL, replacing or swapping the first positional embedding vector $\bm{p}_{1}$ with another position $\bm{p}_{t\neq 1}$ significantly impact the amplitude and position of attention sink.
Replaced position $t$ no 5 10 15 20 25 25
$\textrm{Sink}_{1}^{\epsilon}$(%) 62.28    0.20    2.36    7.73 10.63 10.97 10.21
$\textrm{Sink}_{t}^{\epsilon}$(%) -    0.00    0.00    0.00    0.00    0.00    0.00
Swapped position $t$ no 5 10 15 20 25 25
$\textrm{Sink}_{1}^{\epsilon}$(%) 62.28    1.44    3.73    6.78    8.95    9.42    9.73
$\textrm{Sink}_{t}^{\epsilon}$(%) - 57.63 54.48 52.81 51.70 51.13 50.22
C.2 Attention sink under different data domains